OpenStudy (anonymous):

x^2+3=9x Can someone show me how to solve this?

OpenStudy (anonymous):

\[x ^{2}+3=9x\]

OpenStudy (shamil98):

move the 9x to the other side. x^2 - 9x + 3 = 0 and now you use the quadratic formula. \[\huge x= \frac{ -b \pm \sqrt {b^2 -4ac} }{ 2a }\]

OpenStudy (shamil98):

ax^2 + bx + c = 0 get your values of your equation from this form.

OpenStudy (anonymous):

I don't really understand

OpenStudy (anonymous):

Once you have it n the form ax^2 +bx + c = 0 then use the formula given above to get x. So x^2 - 9x + 3 = 0 has a=1, b=-9, c=3 x = [9 +- sqrt(81-12)/2 = (9 + 8.3)/2 = 8.65 and [9 -8.3]/2 = 0.35 both values of x should solve the equation check x=.35, and the equation = 0 [within rounding errors] x=8.65 also gives 0 w/in rounding errors.: